OpenStudy (anonymous):

Fan and medal c: Check my work with probability? A bag contains 3 red, 5 brown, and 4 yellow potatoes. What is the probability of randomly choosing 2 yellow potatoes without replacement? Write your answer as a decimal. Round to the nearest hundredth if needed. (My work below)

OpenStudy (anonymous):

\[\frac{ 4}{ 12 } \times \frac{ 2 }{ 10 } = \frac{ 1 }{ 15 } = 0.06666667 = 0.07\]